Solve algebraically for x:
3 -(x+10) 1
__= ______ + __
4 4x 2x

Answer by checkley79(3341)

3 -(x+10) 1
__= ______ + __
4 4x 2x
3/4=-(x+10)/4x+1/2x
3/4=(-x-10+1*2)/4x
3/4=(-x-10+2)/4x
3/4=(-x-8)/4x cross multiply.
3*4x=4(-x-8)
12x=-4x-32
12x+4x=-32
16x=-32
x=-32/16
x=-2 ans.